Kent drives his Mazda 150 miles in the same time that it takes Dave to drive his Nissan 135 miles. If Kent averages 5 miles per hour faster than Dave, find their rates.
Answer by mananth(12270)

Dave x mph
Kent 5 faster = x+ 5

Dave 135 miles
Kent 150 miles
Dave time = Kent time
t=d/r
135 / x = 150 /( x + 5 )
cross multiply
135 ( x + 5 ) = 150 x
135 x + 675 = 150 x
150 x - 135 x = 675
15 x = 675
/ 15
x = 45 mph Dave

50 mph Kent
m.ananth@hotmail.ca